I'm doing an online math course and one of the questions says Complete the general solution to y = cos-1(1) for K a whole number. It is wanting me to write out an equation.
Answer by Alan3354(69443)

Complete the general solution to y = cos-1(1) for K a whole number.
-----------
acos(1) = 0 + n*2pi, n = 0,1,2,3...
There's no K involved.